OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (oaktree):

So the equation for a finite arithmetic sum is \[S_{n} =\frac{ n }{ 2 }(a_1 + a_n)\]All we need to do is find n. So we have \[a_n = a_1 + (n-1)d\]And since a_1 is -10 and a_n is 110, and d is obviously 8, we have \[n=\frac{ 110-(-10) }{ 8 } + 1\] Which is 16. So our sum is\[\frac{ 16 }{ 2 }(-10 + 110) = 8(100) = 800\]Good?

OpenStudy (ranga):

Write the sum using summation notation: \[\Large \sum_{n = 1}^{16}(8n - 18)\]

OpenStudy (oaktree):

Oh, summation notation. Sorry. That's right.

OpenStudy (anonymous):

Wait how did you get all thst

OpenStudy (oaktree):

All what?

OpenStudy (anonymous):

no what he wrote

OpenStudy (ranga):

I was basing my answer on OakTree's calculations above.

OpenStudy (oaktree):

I think she means explain where the argument of the summation came from.

OpenStudy (ranga):

We need to find the nth term of the arithmetic series.

OpenStudy (ranga):

This is an arithmetic series with first term -10 and common difference 8. The nth term is given by a_n = a_1 + (n-1)d = -10 + (n-1)(8) = -10 + 8n - 8 a_n = 8n - 18. We need to sum this as n runs from 1 to 16 (16 was determined to be the number of terms in the series by OakTree above).

OpenStudy (ranga):

You can factor the 2 out of the summation.